# AgentTTS: Large Language Model Agent for Test-time Compute-optimal Scaling Strategy in Complex Tasks

### v1

@@ A.6 More Preliminary Experimental Results
A.7 More Main Experimental Results
Fig. 15 is the performance trajectories on WebQSP, Taskbench, and ChatDev datasets.
A.8 Detailed Cases for Interpretability
To illustrate the interpretability benefit from integrating prior insights into the LLM agent for test-time compute allocation, we present three case studies on 2WikiMultiHopQA in Fig. 16. The first row shows Insight 1 enables subtask-specific model use, favoring large models for retrieval and small ones for QA, to guide efficient trials. The second row shows Insight 2 pinpoints the optimal sampling range (5–50), narrowing the search space. The third row shows Insight 3 promotes balanced allocation, favoring one-sample “high-capacity” retrieval and “lower-cost” QA configuration, enabling effective trade-offs for better overall performance.
@@ A.8 Detailed Cases for Interpretability
A.9 Prompt Design in AgentTTS for Search Guidelines and Trial Generation
@@
A.10 Baselines
We compare two baseline categories: LLM-based approaches (AgentHPO Liu et al. [2024a], MLCopilot Zhang et al. [2024a], and LLM_ZS) and traditional methods (Random Search and Bayesian Optimization [Shahriari et al., 2015]).
AgentHPO [Liu et al., 2024a]
targets Hyperparameter Optimization (HPO), addressing limitations of traditional AutoML methods such as high trial cost, complex setup, and limited interpretability by leveraging LLM-powered autonomous agents. It comprises two modules: the Creator, which converts natural language inputs (e.g., dataset, model, and goals) into initial hyperparameters, and the Executor, which trains models, logs outcomes, and returns feedback. The Creator iteratively refines HPs using this feedback, forming a closed-loop optimization. To adapt AgentHPO for our test-time scaling budget allocation setting, we provide AgentHPO with structured inputs similar to our method, including task descriptions, model choices, and total budget. However, AgentHPO assumes LLMs possess sufficient ML knowledge for HPO tasks, an assumption that fails in multi-stage test-time scaling, where domain-specific knowledge is lacking Wang et al. [2024a]. Our method addresses this gap by injecting targeted insights about test-time scaling into the agent’s reasoning process, improving both search efficiency and decision interpretability.
MLCopilot [Zhang et al., 2024a]
focuses on generating machine learning solutions, addressing limitations in traditional AutoML methods, such as Bayesian optimization [Shahriari et al., 2015], which often lack interpretability and struggle to incorporate human prior knowledge, e.g., regarding model architectures. Inspired by human design patterns, where individuals understand a task by recalling relevant experiences or knowledge from prior work or textbooks, MLCopilot emulates this process by leveraging LLMs to suggest solutions for new ML tasks based on historical cases. To adapt MLCopilot to our problem setting, we provide the search history of a similar task to guide the LLM in generating both prior experience and candidate trials iteratively for the target multi-stage task. However, while MLCopilot can extract surface-level patterns from related tasks, it lacks the ability to derive tailored insights necessary for efficient budget allocation in test-time scaling. This limits its effectiveness in the test-time scaling compute-optimal budget allocation problem, where fine-grained domain understanding is essential for optimal resource allocation.
@@ A.10 Baselines
A.11 Detailed Related Work
A.11.1 Test-time Scaling and Compute-optimal Strategy
@@
Inspired by the human reasoning process that tends to engage in deeper, more deliberate thinking, several studies have proposed allocating additional compute during inference to enhance task performance [Wei et al., 2022, Wang et al., 2023]. Further, other works [Brown et al., 2024, Wu et al., 2025] have observed that increasing inference-time compute exhibits a similar pattern akin to training scaling laws: additional inference compute consistently leads to improved task performance. This phenomenon is commonly referred to as test-time scaling (TTS). Test-time scaling technologies can be mainly classified into two categories: sequential scaling and parallel scaling. In sequential scaling, it enhances test-time computation by generating progressively longer solutions along the sequence dimension. The most prevalent method is self-revision, where first generate an initial response and then iteratively evaluate and refine it based on self-assessment [Madaan et al., 2023, Gou et al., 2024, Snell et al., 2025]. Because sequential scaling relies on the model’s ability to generate reasonably accurate initial responses, it is generally more effective for simpler tasks [Snell et al., 2025]. On the other hand, parallel scaling generates multiple independent solutions in parallel and aggregates them into a final answer. Common solution-level aggregation methods are Best-of-N [Sun et al., 2024, Gui et al., 2024], which sample N complete solutions and then select the best one according to a verifier. While Tree-Search algorithms, viewed as parallel scaling at the token or step level, are against a process-based reward model to search top-K intermediate steps and further explore, including Beam Search [Yu et al., 2024, Xie et al., 2023] and Monte Carlo Tree Search (MCTS) [Wu et al., 2025, Snell et al., 2025, Hao et al., 2023, Wan et al., 2024, Chen et al., 2024, Zhang et al., 2023b]. However, they typically rely on explicit guidance signals for candidate selection. An alternative line of work directly employs an LLM as the fuser to aggregate candidate solutions, offering greater generalization and flexibility [Jiang et al., 2023, Li et al., 2025, Saad-Falcon et al., 2024]. Zeng et al. [2025] argues that parallel scaling provides better coverage and scalability than sequential scaling, as extended sequential reasoning chains may corrupt intermediate correct outputs. Similarly, Snell et al. [2025] demonstrates that parallel scaling is more suitable for complex tasks, as it requires only the ability to generate correct final answers. Therefore, we adopt a parallel test-time scaling strategy—specifically, repeated sampling with fusion, to address complex multi-stage reasoning tasks, leveraging the strengths of small models in diverse sampling and robust aggregation.
@@
The configuration of test-time scaling strategies plays a critical role in performance improvement, giving rise to the broader problem of optimally allocating compute and budget, commonly referred to as the test-time compute-optimal scaling strategy. Recent studies [Brown et al., 2024, Wu et al., 2025, Liu et al., 2025a, Yue et al., 2025, Snell et al., 2025, Wang et al., 2024b] have shown that, in certain settings, smaller models can outperform large language models under the same compute budget. These works explore how to choose between training scaling and test-time scaling, as well as how to select among different test-time scaling techniques to optimize performance. For instance, Snell et al. [2025] demonstrate that task difficulty strongly influences the optimal scaling strategy: for moderately difficult tasks, parallel search with small models is preferred, while for simpler tasks, sequential scaling with large models is more effective. Based on this observation, they propose a difficulty predictor to guide scaling decisions dynamically. Liu et al. [2025a] further extend this line of work by showing that the choice of optimal TTS strategies is highly sensitive to the design of reward functions. Yue et al. [2025] propose using a linear model to fit key factors influencing strategy selection within retrieval-augmented generation (RAG) systems. Wu et al. [2025] introduce Reward Balanced Search (REBASE), a novel tree search algorithm that combines weighted voting to achieve a Pareto-optimal trade-off between accuracy and inference cost. However, these approaches primarily focus on single-stage task scenarios and do not systematically address the challenge of budget allocation across subtasks in multi-stage complex tasks. In this work, we extend the test-time compute-optimal scaling framework to multi-stage complex tasks.
A.11.2 LLM for Hyperparameter Optimization
@@
LLMs have emerged as promising tools for enhancing the efficiency and effectiveness of hyperparameter optimization (HPO) by leveraging contextual understanding and prior knowledge [Gu et al., 2024, Wang et al., 2024c, Zhang et al., 2025, Liu et al., 2025b, Wang et al., 2024a], compared to traditional Automated Machine Learning (AutoML) approaches such as Bayesian Optimization (BO) [Shahriari et al., 2015]. Existing LLM-based HPO studies can be mainly categorized into two directions: (1) using LLMs to reduce the search space for traditional methods, and (2) enabling LLMs to autonomously search optimal hyperparameters. On one hand, LLMs can significantly reduce the vast search spaces in Neural Architecture Search (NAS) and HPO [Yu et al., 2023, Morris et al., 2024, Luo et al., 2024, Liu et al., 2024a]. For example, GPT-NAS [Yu et al., 2023] combines GPT models with evolutionary algorithms [Zhang et al., 2021] to rapidly prune low-quality architectures, enhancing search efficiency. Morris et al. [2024] introduce "Evolution of Thought," allowing LLMs to refine architectures and hyperparameters based on feedback iteratively. AutoM3L [Luo et al., 2024] integrates external tools such as Ray.Tune for hyperparameter tuning, with LLMs recommending optimized search spaces. Llambo [Liu et al., 2024a] formulates BO problems in natural language, enabling LLMs to propose high-potential candidates and iteratively optimize based on historical observations.
@@
On the other hand, several works [Chen et al., 2023, Lin et al., 2025, Zheng et al., 2023, Zhang et al., 2023a, Achiam et al., 2023, Hollmann et al., 2023, Zhang et al., 2024a, b] empower LLMs to search hyperparameter configurations autonomously. AutoMMLab [Yang et al., 2025b] and GENIUS [Zheng et al., 2023] treat LLMs as black-box optimizers, iteratively refining configurations based on evaluation feedback. CAAFE [Hollmann et al., 2023] focuses on tabular data, where LLMs generate new semantically meaningful features based on dataset descriptions. Moreover, LLM-based agent frameworks have emerged, leveraging feedback from machine learning platforms or historical experiments. For instance, AutoML-GPT [Zhang et al., 2023a] automates the pipeline from preprocessing to model selection and hyperparameter tuning, adjusting strategies based on experimental logs. MLCopilot [Zhang et al., 2024a] predicts hyperparameters for unseen tasks by canonicalizing prior task experiences and interactively refining solutions through LLM reasoning. AgentHPO [Liu et al., 2024a] autonomously processes task descriptions, conducts experiments, and iteratively improves hyperparameter search based on accumulated trials. Building upon these advancements, we extend the LLM-agent framework to the domain of test-time scaling budget allocation in multi-stage complex tasks, aiming to search for compute-optimal scaling configurations.
A.12 Limitations

### v2

@@ A.6 More Preliminary Experimental Results
A.7 Prompt Design in AgentTTS for Search Guidelines and Trial Generation
@@
A.8 Detailed Cases for Interpretability
A.9 Baselines
We compare two baseline categories: LLM-based approaches (AgentHPO [20], MLCopilot [56], and LLM_ZS) and traditional methods (Random Search and Bayesian Optimization [31]).
AgentHPO [20]
targets Hyperparameter Optimization (HPO), addressing limitations of traditional AutoML methods such as high trial cost, complex setup, and limited interpretability by leveraging LLM-powered autonomous agents. It comprises two modules: the Creator, which converts natural language inputs (e.g., dataset, model, and goals) into initial hyperparameters, and the Executor, which trains models, logs outcomes, and returns feedback. The Creator iteratively refines HPs using this feedback, forming a closed-loop optimization. To adapt AgentHPO for our test-time scaling budget allocation setting, we provide AgentHPO with structured inputs similar to our method, including task descriptions, model choices, and total budget. However, AgentHPO assumes LLMs possess sufficient ML knowledge for HPO tasks, an assumption that fails in multi-stage test-time scaling, where domain-specific knowledge is lacking. Our method addresses this gap by injecting targeted insights about test-time scaling into the agent’s reasoning process, improving both search efficiency and decision interpretability.
MLCopilot [56]
focuses on generating machine learning solutions, addressing limitations in traditional AutoML methods, such as Bayesian optimization [31], which often lack interpretability and struggle to incorporate human prior knowledge, e.g., regarding model architectures. Inspired by human design patterns, where individuals understand a task by recalling relevant experiences or knowledge from prior work or textbooks, MLCopilot emulates this process by leveraging LLMs to suggest solutions for new ML tasks based on historical cases. To adapt MLCopilot to our problem setting, we provide the search history of a similar task to guide the LLM in generating both prior experience and candidate trials iteratively for the target multi-stage task. However, while MLCopilot can extract surface-level patterns from related tasks, it lacks the ability to derive tailored insights necessary for efficient budget allocation in test-time scaling. This limits its effectiveness in the test-time scaling compute-optimal budget allocation problem, where fine-grained domain understanding is essential for optimal resource allocation.
@@ A.9 Baselines
A.10 Detailed Related Work
A.10.1 Test-time Scaling and Compute-optimal Strategy
@@
Inspired by the human reasoning process that tends to engage in deeper, more deliberate thinking, several studies have proposed allocating additional compute during inference to enhance task performance [44, 42]. Further, other works [2, 45] have observed that increasing inference-time compute exhibits a similar pattern akin to training scaling laws: additional inference compute consistently leads to improved task performance. This phenomenon is commonly referred to as test-time scaling (TTS). Test-time scaling technologies can be mainly classified into two categories: sequential scaling and parallel scaling. In sequential scaling, it enhances test-time computation by generating progressively longer solutions along the sequence dimension. The most prevalent method is self-revision, where first generate an initial response and then iteratively evaluate and refine it based on self-assessment [24, 6, 33]. Because sequential scaling relies on the model’s ability to generate reasonably accurate initial responses, it is generally more effective for simpler tasks [33]. On the other hand, parallel scaling generates multiple independent solutions in parallel and aggregates them into a final answer. Common solution-level aggregation methods are Best-of-N [35, 9], which sample N complete solutions and then select the best one according to a verifier. While Tree-Search algorithms, viewed as parallel scaling at the token or step level, are against a process-based reward model to search top-K intermediate steps and further explore, including Beam Search [53, 46] and Monte Carlo Tree Search (MCTS) [45, 33, 10, 37, 5, 59]. However, they typically rely on explicit guidance signals for candidate selection. An alternative line of work directly employs an LLM as the fuser to aggregate candidate solutions, offering greater generalization and flexibility [13, 17, 30]. Zeng et al. [55] argues that parallel scaling provides better coverage and scalability than sequential scaling, as extended sequential reasoning chains may corrupt intermediate correct outputs. Similarly, Snell et al. [33] demonstrates that parallel scaling is more suitable for complex tasks, as it requires only the ability to generate correct final answers. Therefore, we adopt a parallel test-time scaling strategy—specifically, repeated sampling with fusion, to address complex multi-stage reasoning tasks, leveraging the strengths of small models in diverse sampling and robust aggregation.
@@
The configuration of test-time scaling strategies plays a critical role in performance improvement, giving rise to the broader problem of optimally allocating compute and budget, commonly referred to as the test-time compute-optimal scaling strategy. Recent studies [2, 45, 19, 54, 33, 41, 40] have shown that, in certain settings, smaller models can outperform large language models under the same compute budget. These works explore how to choose between training scaling and test-time scaling, as well as how to select among different test-time scaling techniques to optimize performance. For instance, Snell et al. [33] demonstrate that task difficulty strongly influences the optimal scaling strategy: for moderately difficult tasks, parallel search with small models is preferred, while for simpler tasks, sequential scaling with large models is more effective. Based on this observation, they propose a difficulty predictor to guide scaling decisions dynamically. Liu et al. [19] further extend this line of work by showing that the choice of optimal TTS strategies is highly sensitive to the design of reward functions. Yue et al. [54] propose using a linear model to fit key factors influencing strategy selection within retrieval-augmented generation (RAG) systems. Wu et al. [45] introduce Reward Balanced Search (REBASE), a novel tree search algorithm that combines weighted voting to achieve a Pareto-optimal trade-off between accuracy and inference cost. However, these approaches primarily focus on single-stage task scenarios and do not systematically address the challenge of budget allocation across subtasks in multi-stage complex tasks. In this work, we extend the test-time compute-optimal scaling framework to multi-stage complex tasks.
A.10.2 LLM for Hyperparameter Optimization
@@
LLMs have emerged as promising tools for enhancing the efficiency and effectiveness of hyperparameter optimization (HPO) by leveraging contextual understanding and prior knowledge [8, 43, 61, 22, 38], compared to traditional Automated Machine Learning (AutoML) approaches such as Bayesian Optimization (BO) [31]. Existing LLM-based HPO studies can be mainly categorized into two directions: (1) using LLMs to reduce the search space for traditional methods, and (2) enabling LLMs to autonomously search optimal hyperparameters. On one hand, LLMs can significantly reduce the vast search spaces in Neural Architecture Search (NAS) and HPO [52, 25, 23, 20]. For example, GPT-NAS [52] combines GPT models with evolutionary algorithms [60] to rapidly prune low-quality architectures, enhancing search efficiency. Morris et al. [25] introduce "Evolution of Thought," allowing LLMs to refine architectures and hyperparameters based on feedback iteratively. AutoM3L [23] integrates external tools such as Ray.Tune for hyperparameter tuning, with LLMs recommending optimized search spaces. Llambo [20] formulates BO problems in natural language, enabling LLMs to propose high-potential candidates and iteratively optimize based on historical observations.
@@
On the other hand, several works [4, 18, 62, 58, 1, 12, 56] empower LLMs to search hyperparameter configurations autonomously. AutoMMLab [49] and GENIUS [62] treat LLMs as black-box optimizers, iteratively refining configurations based on evaluation feedback. CAAFE [12] focuses on tabular data, where LLMs generate new semantically meaningful features based on dataset descriptions. Moreover, LLM-based agent frameworks have emerged, leveraging feedback from machine learning platforms or historical experiments. For instance, AutoML-GPT [58] automates the pipeline from preprocessing to model selection and hyperparameter tuning, adjusting strategies based on experimental logs. MLCopilot [56] predicts hyperparameters for unseen tasks by canonicalizing prior task experiences and interactively refining solutions through LLM reasoning. AgentHPO [20] autonomously processes task descriptions, conducts experiments, and iteratively improves hyperparameter search based on accumulated trials. Building upon these advancements, we extend the LLM-agent framework to the domain of test-time scaling budget allocation in multi-stage complex tasks, aiming to search for compute-optimal scaling configurations.
A.11 Limitations